# PIS3R: Very Large Parallax Image Stitching via Deep 3D Reconstruction

### v2

@@ III-C Image Stitching via Point Cloud Reprojection
Subsequently, we construct a pixel grid $\mathcal{P}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times 4}$, whose spatial resolution $(H_{i},W_{i})$ is dynamically determined from the coordinate range of the projected points $P_{uv}^{i}$. For each integer pixel location $(x,y)$, we define the set of 3D points that project to it as:
@@
Among these, we select the point with minimum depth in the camera coordinate system:
@@
$\mathcal{P}_{i}(x,y)=\mathcal{G}\big(c_{k},\,Z_{c}^{i}(k)\big),$
where $P_{w}(j)$ denotes the $j$-th 3D point in world coordinates, $Z_{c}^{i}(j)$ its depth in the $i$-th camera frame, and $P_{uv}^{i}(j)$ its 2D projection onto the corresponding image plane in pixel coordinates.
To alleviate the burden on the diffusion model during refinement, we further process the holes within the projected regions. Let $S_{i}$ denote the set of all filled pixel positions. We then compute the convex hull of $S_{i}$, denoted $\mathcal{H}_{i}$, which represents the minimal convex region that encompasses all reprojected pixels:
@@
$\mathcal{H}_{i}=\mathrm{Convex\ Hull}(S_{i}).$ (7)
Based on $S_{i}$ and $\mathcal{H}_{i}$, we construct three binary masks on the integer lattice:
$M_{\mathrm{f}}(x,y)$ $=\mathbf{1}\!\big[(x,y)\in S_{i}\big],$ (8)
$M_{\mathrm{r}}(x,y)$ $=\mathbf{1}\!\big[(x,y)\in\mathcal{H}_{i}\big],$
$M_{\mathrm{h}}(x,y)$ $=M_{\mathrm{r}}(x,y)\big(1-M_{\mathrm{f}}(x,y)\big),$
where, $M_{\mathrm{f}}$ indicates the originally observed (filled) pixels, $M_{\mathrm{r}}$ covers all integer pixels lying inside or on the boundary of the convex hull, and $M_{\mathrm{h}}$ identifies interior hole pixels—those within $\mathcal{H}_{i}$ but not in $S_{i}$.
For any hole pixel $(x,y)$ with $M_{\mathrm{h}}(x,y)=1$, we define a neighborhood $\mathcal{N}_{K}(x,y)$ consisting of the $K$ nearest filled pixels within the hull:
$\mathcal{N}_{K}(x,y)\subseteq\left\{(u,v)\,\middle|\,M_{\mathrm{f}}(u,v)=1,\ M_{\mathrm{r}}(u,v)=1\right\}.$ (9)
Each neighbor is assigned a Gaussian weight based on its Euclidean distance to $(x,y)$:
$\alpha\!\big((x,y),(u,v)\big)=\exp\!\left(-\frac{\|(x-u,\ y-v)\|_{2}^{2}}{2\sigma^{2}}\right).$ (10)
The hole color $\mathbf{H}_{i}(x,y)$ is then computed as a distance-weighted average of the RGB values of these neighbors:
$\mathbf{H}_{i}(x,y)=\frac{\sum_{(u,v)\in\mathcal{N}_{K}(x,y)}\alpha\!\big((x,y),(u,v)\big)\,\mathbf{C}_{i}(u,v)}{\sum_{(u,v)\in\mathcal{N}_{K}(x,y)}\alpha\!\big((x,y),(u,v)\big)},$ (11)
where $\mathbf{C}_{i}(u,v)\in\mathbb{R}^{3}$ denotes the RGB color corresponding to the 3D point.
Finally, the complete reprojected image $\hat{S}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times 3}$ is obtained by rasterizing the pixel grid as follows:
$\hat{S}_{i}(x,y)=\begin{cases}\mathbf{C}_{i}(x,y),&\text{if }M_{\mathrm{f}}(x,y)=1,\\ \mathbf{H}_{i}(x,y),&\text{if }M_{\mathrm{h}}(x,y)=1,\\ \mathbf{b},&\text{otherwise},\end{cases}$ (12)
where $\mathbf{b}\in\mathbb{R}^{3}$ is a predefined background color (e.g., black). This formulation ensures that observed pixels retain their original appearance, interior holes are smoothly inpainted from spatially proximate valid samples, and all remaining regions are filled with the background.

### v3

@@ III-C Image Stitching via Point Cloud Reprojection
Subsequently, we construct a pixel grid $\mathcal{P}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times 4}$, whose spatial resolution $(H_{i},W_{i})$ is dynamically determined from the coordinate range of the projected points $P_{uv}^{i}$. For each integer pixel location $(x,y)$, we define the set of 3D points that project to it as Eq. 4, then we select the point with minimum depth in the camera coordinate system as Eq. 5:
@@
$\mathcal{P}_{i}(x,y)=[C_{k},\,Z_{c}^{i}(k)],$
where $P_{w}(j)$ denotes the $j$-th 3D point in world coordinates, $Z_{c}^{i}(j)$ is the depth of $P_{w}(j)$ in the $i$-th camera coordinate, and $P_{uv}^{i}(j)$ its 2D projection onto the corresponding image plane in pixel coordinates, $C_{j}$ is the RGB of $P_{w}(j)$.
To alleviate the artifacts caused by sparse point projection while preserving valid structural voids (e.g., occlusion gaps), we propose a density-aware refinement strategy based on morphological operations. Let $S_{i}$ denote the set of discrete pixel coordinates obtained from the projection:
@@
Then, we generate a valid rendering mask using morphological closing. This operation bridges small gaps between sparsely projected points without altering the global topology. We first define a binary occupancy mask $M_{\mathrm{occ}}$ indicating the raw projected pixels:
$M_{\mathrm{occ}}(x,y)=\mathbf{1}\!\big[(x,y)\in S_{i}\big].$ (7)
The geometrically closed mask, denoted as $M_{\mathrm{closed}}$, is derived by applying a morphological dilation followed by an erosion with a structuring element $B$ (e.g., a disk kernel):
$M_{\mathrm{closed}}=(M_{\mathrm{occ}}\oplus B)\ominus B,$ (8)
where $\oplus$ and $\ominus$ denote the dilation and erosion operators, respectively. We then identify the set of interior hole pixels, denoted by the mask $M_{\mathrm{hole}}$, as the difference between the morphologically closed region and the original projection:
$M_{\mathrm{hole}}(x,y)=M_{\mathrm{closed}}(x,y)\big(1-M_{\mathrm{occ}}(x,y)\big).$ (9)
For any hole pixel $(x,y)$ where $M_{\mathrm{hole}}(x,y)=1$, we apply an inpainting algorithm based on the Fast Marching Method (FMM)  [telea2004image]. This method propagates color information from the boundary of the occupied region into the holes, effectively maintaining sharp edges and local consistency.
Finally, the refined reprojected image $\hat{S}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times 3}$ is composed by merging the original projection, the inpainted regions, and the background:
$\hat{S}_{i}(x,y)=\begin{cases}\mathbf{C}_{i}(x,y),&\text{if }M_{\mathrm{occ}}(x,y)=1,\\ \mathbf{H}_{i}(x,y),&\text{if }M_{\mathrm{hole}}(x,y)=1,\\ \mathbf{b},&\text{otherwise},\end{cases}$ (10)
where $\mathbf{b}\in\mathbb{R}^{3}$ is a predefined background color. This formulation ensures that artifacts arising from projection sparsity are seamlessly filled, while legitimate structural voids and background regions remain preserved.